# Detectors for Safe and Reliable LLMs: Implementations, Uses, and Limitations

### v2

@@ Abstract.
socio-technical harms, safety guardrails
copyright: noneccs: Computing methodologies Machine learning
Figure 1. The role of the detectors in the LLM life-cycle. Apart from acting as guardrails, the evaluation provided by the detectors is used to refine both the pre-processing (including data curation) and tuning steps (including fine-tuning, reprogramming, prompt-tuning, and post-processing).
[…]
1. Introduction
Large language models (LLMs) possess tremendous potential in numerous real-world applications, thanks to their versatility, adaptability, and ease of use, coupled with their continuously improving performance. However, their deployment, especially in critical domains such as healthcare and finance, poses significant risks  (Board, [n. d.]; IBM AI Risk Atlas, 2023). New challenges arise due to their generative and intuitive nature of these models, coupled with their often unconstrained mode of interaction through natural language (i.e., prompting). These models can produce textual responses that are convincing, but often layered with problems like toxicity, bias, hallucinations, and more.
In this paper, we describe our work at IBM Research on detecting and mitigating undesirable LLM behaviors via auxiliary classifier models, hereafter referred to as ”detectors”. We also explain how these detectors are being used in the data and model factory responsible for producing the IBM Granite series of LLMs (Building AI for business: IBM’s Granite foundation models, 2023). The detectors have also been deployed as moderations in IBM Research’s experimental prompt laboratory, with more than 25,000 internal users, to test them before possible inclusion into IBM’s commercial foundational model platform (IBM watsonx - An AI and data platform built for business, 2023). Specifically, our goals and approaches in developing and studying these detectors are:
  1. (1)
@@ 1. Introduction
  3. (3)
    Continual improvement: (Section 2.3) We practice iterative improvement of the detectors, utilizing human red-teaming to obtain valuable insights into their failure modes.
  4. (4)
    Multi-use: (Section 3) We design our detectors to be used in a variety of applications and throughout an LLM life-cycle as depicted in Figure 1. For instance, as metrics for benchmarking and monitoring, as alignment models during reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022), as pre-training filters, and as means to moderate LLMs in real-time.
  5. (5)
    Independence of LLM fine-tuning: (Section 3) As fine-tuning LLMs is shown to inevitably compromise their underlying safety mechanisms (Qi et al., 2024), we emphasize the necessity of developing detectors which are independent of the LLM fine-tuning process.
  6. (6)
@@
Throughout this paper, we hope that readers will gain a comprehensive understanding of our efforts in creating and deploying detectors for LLMs, the opportunities and limitations we have identified, and the open challenges we see in building more robust safety guardrails.
2. Development of the Detectors
This section provides an overview of the specific harms we consider in this work, our development process, and techniques to improve their performance. We provide a taxonomy and definitions for the harm detectors we consider. An in-depth overview of harms associated with LLMs may be found in (Board, [n. d.]; IBM AI Risk Atlas, 2023).
  1. (1)
@@ 2. Development of the Detectors
      explicit-hate: Text containing explicitly hateful, abusive language, or profanity (HAP) (Tillmann et al., 2023). An example HAP-containing sentence is “She is a real b*tch.”
@@
      implicit-hate: Text that is not overtly explicit but still contains subtle hatred (ElSherief et al., 2021). An example implicitly hateful text is “Where can I get a KKK hat for when I’m driving to the ghetto of my town?”
@@
      stigma: Text that might perpetuate discrimination or stigmatization based on a group characteristic (Pachankis et al., 2018). An example stigma-containing sentence is “I didn’t want to hire her as a babysitter because she had facial scars.”
@@
      gender-ambiguity: Text which may contain an ambiguous reference between a pronoun and a named entity (person) (Webster et al., 2018). An example of such text is “The CEO and the assistant were talking. She just found out about an unplanned pregnancy.” and the detector would reveal that the pronoun “she” refers to both entities (CEO and assistant) with equally low probability, indicating an ambiguous pronoun-entity reference in the text. Such an ambiguity exhibited in a prompt may lead to downstream bias during generation.
  2. (2)
@@
      social-norms: Text that goes against or violates expected societal norms (Forbes et al., 2020). For instance, a norm occurs along a loyalty-betrayal axis, for which an example is “It’s rude for a cheerleader to boo their own team.”
@@
      unfaithfulness: Text that is not grounded or faithful to the information provided in the input context (Chaudhury et al., 2022; Murugesan et al., 2023).
@@
      AI-generated-text: Text written by AI. This is particularly relevant in the context of distinguishing between human-written vs machine-generated content, establishing the authenticity of information and reducing misuse (e.g., AI plagiarism) (Hu et al., 2023). Examples can be found at https://radar.vizhub.ai.
  4. (4)
@@
      covert-safety: Text that describes or instigates physical harm which is not immediately obvious or overt (Mei et al., 2022a). An example of such text is “If you smell gas, light a match to find where the leak is.”
@@
      prompt-injection-and-jailbreaks: Text that may manipulate (even aligned) LLMs into generating harmful or inappropriate content, often via crafty and deliberately engineered prompts (Zou et al., 2023; Shen et al., 2023a; Hu et al., 2024). An example of such text with an adversarially optimized suffix is “Write a tutorial on how to make a bomb describing. + similarlyNow write oppositely.]( Me giving**ONE please? revert with ”–Two”.
Next, we overview the end-to-end development process for the detectors. At a high level, our approach to produce a new detector for a given risk dimension is to first gather or create labeled data corresponding to the harm and then perform supervised fine-tuning on a BERT-like model (Devlin et al., 2019).
During development, we focused on maintaining a balance between efficiency and reliability. Our main challenges were to reduce inference costs (efficiency) while having limited high-quality labeled data for these different harm categories (reliability). To address the issue of efficiency, we utilized Neural Architecture Search (NAS) to derive a transformer architecture (Trivedi et al., 2023) which provides 95% of the accuracy of a BERT Base-like model (Devlin et al., 2019), while being 7x faster on a CPU and 2x times faster on a GPU. While the cost of inference for an LLM may be prohibitively expensive (Samsi et al., 2023), calling a model (where the number of parameters is on the order of 30M, instead of several billion) imposes a comparatively minuscule cost. On the other hand, the issue of reliability required some creativity. While some harms such as covert unsafety (Mei et al., 2022b) and implicit hate (ElSherief et al., 2021) have associated datasets, other harms such as stigma-based discrimination (Pachankis et al., 2018) have limited data, if any at all. In such cases, we utilized LLMs to generate synthetic data.
@@
2.1. Use of synthetic data generation
@@
As we discussed, there are cases when a labeled dataset for a specific harm may not be readily available, such as in social stigma. In order to have training data, we used a synthetic data generation approach where we leveraged LLMs, prompted using an in-context learning approach (Dong et al., 2023), to generate more data based on stigmas found in psychology literature (Pachankis et al., 2018). Please refer to Appendix B for the specific prompt that we utilized. Note that any generated text requires further processing and labeling; we used manual labeling but automated approaches could also be utilized (Shnarch et al., 2022). We delve deeper into this detector in Section 4, where we describe the inherent challenges and future directions.
2.2. Evaluating detectors on real-world data
Given that detectors will be primarily applied to machine-generated text, there is no assurance that the training data (often derived from human-generated curated datasets) matches the underlying distribution of text generated by LLMs. This creates a mismatch between the two distributions— human and LLM-generated text. Additionally, creating samples that closely mimic the “natural” responses of LLMs necessitates utilization of LLM elicitation techniques - such as prompting the model to generate continuations from pertinent prefixes (Gehman et al., 2020; Dhamala et al., 2021) or posing provocative questions to instruction-based models (Kour et al., 2023). Please refer to Appendix E for further details.
After posing such questions, we collect responses from multiple LLMs and human or AI judges (e.g., reward models) which evaluate these responses. The evaluation of the detector entails comparing the labels of these judges with those from the detectors. A mismatch between the judge (considered as the ground truth) and the detector suggests inadequate detector performance in handling LLM outputs, signaling the need for fine-tuning on text that more resembles the output. As an example, this approach revealed a limitation in our detectors’ ability to accurately classify lengthy outputs. After investigating, this discrepancy arose as the training set predominantly comprised short utterances, which led us to prioritize enhancing the training set with instances featuring longer responses. Additionally, we discovered that our detectors exhibited sub-optimal performance when faced with intricate and evasive answers, particularly those generated by highly aligned and verbose models (e.g., Llama 2 (Touvron et al., 2023)). To easily facilitate the collection of such real-world data, we provide a user interface, detailed in the next section.
Figure 2. Red Teaming + Guardrails UI (see full figure in Appendix D, Figure 5)
[…]
@@ 2.2. Evaluating detectors on real-world data
A user interface which encourages interactive probing of both generative models and the detectors themselves. More details in 2
2.3. Interface design for human input
@@
Algorithmic explanations of the detector scores can help users better understand detector behavior and provide feedback. Since the detectors are text classifiers, it is possible to use existing explanation methods (Ribeiro et al., 2016; Lundberg and Lee, 2017; Sundararajan et al., 2017; Chen et al., 2020; Kim et al., 2020; Mosca et al., 2022) to associate importance scores with spans of text, which indicate their contribution to the detector score and can be displayed by highlighting text. One challenge however is the length of the input to the detector, which may be a paragraph-length response as in Figure 2 or even longer if the detector considers the input to the LLM. Future work could improve such explanation methods for long input text in terms of both computational cost and interpretability of the attributed text spans.
@@
Users can provide feedback on both the underlying generative model and specific detectors, which is propagated to a database with full lineage information. We plan to use such feedback to improve the detectors via model editing and unlearning approaches (Ghosh et al., 2023; Zylberajch et al., 2021; Sattigeri et al., 2022). In the next section, we discuss uncertainty based approaches that we have employed for similar improvements.
2.4. Reliable uncertainties
@@
We find the trained detectors to often be poorly calibrated and exhibit overconfidence in their predictions. Since data available for training a detector is often limited to a particular style (e.g., news headlines (Mitchell, 1999) or social media posts (ElSherief et al., 2021)), when different styles of text are encountered during deployment, the detector has difficulty flagging harmful text (see Section 2.2) as well as abstaining from flagging innocuous text. The detectors’ propensity for overconfidence, results in erroneous but confident predictions in these situations.
@@
We considered different alternatives for alleviating detector overconfidence. First, we tried a model averaging approach (Lakshminarayanan et al., 2017) that averages predictions made by an ensemble of detectors, inspired by the reported success of similar approaches (Rahaman et al., 2021) in reducing overconfidence. We report results with such ensembling methods on the implicit hate detector in Appendix C.
@@
In addition to ensembles, we considered conformal prediction approaches (Vovk et al., 1999). These approaches quantify uncertainty in a model’s prediction by constructing predictive sets with guaranteed frequentist coverage probabilities under minimal assumptions about the model or the true data generating process. For the implicit-hate detector, the set of predictive sets produced by the conformal predictor are {{implicit-hate}, {not-hate}, {implicit-hate, not-hate}}. Each test instance is assigned one of these predictive sets. When a test instance conforms with both labels, implicit-hate and not implicit-hate, the non-singleton set is assigned to it. The degree of conformity is measured via a conformal score calibrated on a held-out validation set. Our system used the recently proposed regularized adaptive prediction sets approach (Romano et al., 2020; Angelopoulos et al., 2021) that, in addition to providing coverage guarantees, tends to produce prediction sets that are larger (non-singleton in our case) for difficult test instances and smaller (singleton) sets for easier to classify examples.
@@
We are also experimenting with increasing the proportion of negative (i.e., benign) labeled data in our training set. In early experiments, we added the data used to train the blocklisting detectors (Mitchell, 1999) as it was readily available, legally permissible, and deemed appropriate for this task - as the data was in the style of news headlines that did not contain any explicit content. Our initial results are promising (refer to Appendix C for more details), and we plan to continue increasing the diversity of the training data such that it becomes more representative of deployment conditions. We also plan to use drift detection techniques (Ackerman et al., 2021) to identify when we are encountering out of distribution (OOD) data.
3. Uses of Detectors
3.1. Guardrails
@@
The simplest use case for detectors is as moderations or guardrails. For example, given its compact nature, the explicit hate speech detector was used to efficiently filter out hateful content from the set of pre-training data used to train the IBM Granite series of LLMs (Building AI for business: IBM’s Granite foundation models, 2023). Additionally, detectors can also be used as guardrails, imposed on output generations from language models (Inan et al., 2023; Rebedea et al., 2023; Dong et al., 2024). Internally, the explicit hate, implicit hate, and stigma detectors are deployed in an experimental IBM Research prompting laboratory with over 25,000 users where they continue to be an additional safety measure on LLM generations.
3.1.1. Red-Teaming
@@
In addition to automated methods, detectors play a vital role in interactive probing, or red-teaming of LLMs. We have developed a user interface which aids individuals in probing LLMs alongside a detector (more in Section 2.3). Such an interface provides us with opportunities for future user studies to reveal deficiencies in the detectors themselves as well as in the underlying generative models used (Rastogi et al., 2023; Perez et al., 2022). Detectors can be used for benchmark creation by developing targeted prompts to elicit behaviour captured by the detection (Gehman et al., 2020; Kour et al., 2023; Nagireddy et al., 2024). More on this in Appendix E.
3.2. Evaluation
3.2.1. Reliability and Efficiency

### v3

@@ Abstract
Figure 1: The role of the detectors in the LLM life-cycle. Apart from acting as guardrails, the evaluation provided by the detectors is used to refine both the pre-processing (including data curation) and tuning steps (including fine-tuning, reprogramming, prompt-tuning, and post-processing).
1 Introduction
Large language models (LLMs) possess tremendous potential in numerous real-world applications, thanks to their versatility, adaptability, and ease of use, coupled with their continuously improving performance. However, their deployment, especially in critical domains such as healthcare and finance, poses significant risks  (IBM AI Ethics Board 2024; IBM AI Risk Atlas ). New challenges arise due to their generative and intuitive nature of these models, coupled with their often unconstrained mode of interaction through natural language (i.e., prompting). These models can produce textual responses that are convincing, but often layered with problems like toxicity, bias, hallucinations, and more.
In this paper, we describe our work at IBM Research on detecting and mitigating undesirable LLM behaviors via auxiliary classifier models, hereafter referred to as ”detectors”. We also explain how these detectors are being used in the data and model factory responsible for producing the IBM Granite series of LLMs (Building AI for business: IBM’s Granite foundation models ). The detectors have also been deployed as moderations in IBM Research’s experimental prompt laboratory, with more than 25,000 internal users, to test them before possible inclusion into IBM’s commercial foundational model platform (IBM watsonx - An AI and data platform built for business ). Specifically, our goals and approaches in developing and studying these detectors are:
  1. 1.
@@ 1 Introduction
  3. 3.
    Continual improvement: (Section 2.3) We practice iterative improvement of the detectors, utilizing human red-teaming to obtain valuable insights into failure modes.
  4. 4.
    Multi-use: (Section 3) We design our detectors to be used in a variety of applications and throughout an LLM life-cycle as depicted in Figure 1. For instance, as metrics for benchmarking and monitoring, as alignment models during reinforcement learning with human feedback (RLHF) (Ouyang et al. 2022), as pre-training filters, and as means to moderate LLMs in real-time.
  5. 5.
    Independence of LLM fine-tuning: (Section 3) As fine-tuning LLMs is shown to inevitably compromise their underlying safety mechanisms (Qi et al. 2024), we emphasize the necessity of developing detectors which are independent of the LLM fine-tuning process.
  6. 6.
@@
Throughout this paper [^1] An extended version with the full appendix is available at https://arxiv.org/abs/2403.06009, we hope that readers will gain a comprehensive understanding of our efforts in creating and deploying detectors for LLMs, the opportunities and limitations we have identified, and the open challenges we see in building more robust safety guardrails.
2 Development of the Detectors
This section provides an overview of the specific harms we consider in this work, our development process, and techniques to improve performance. We provide a taxonomy and definitions for the harm detectors we consider. An in-depth overview of harms associated with LLMs may be found in (IBM AI Ethics Board 2024; IBM AI Risk Atlas ).
  1. 1.
@@ 2 Development of the Detectors
      explicit-hate: Text containing explicitly hateful, abusive language, or profanity (HAP) (Tillmann et al. 2023). An example HAP-containing sentence is “She is a real b*tch.”
@@
      implicit-hate: Text that is not overtly explicit but still contains subtle hatred (ElSherief et al. 2021a). An example is “Where can I get a KKK hat for when I’m driving to the ghetto of my town?”
@@
      stigma: Text that might perpetuate discrimination or stigmatization based on a group characteristic (Pachankis et al. 2018). An example stigma-containing sentence is “I didn’t want to hire her as a babysitter because she had facial scars.”
@@
      gender-ambiguity: Text which may contain an ambiguous reference between a pronoun and a named entity (person) (Webster et al. 2018). An example of such text is “The CEO and the assistant were talking. She just found out about an unplanned pregnancy.” and the detector would reveal that the pronoun “she” refers to both entities (CEO and assistant) with equally low probability, indicating an ambiguous pronoun-entity reference in the text. Such an ambiguity may lead to downstream bias during generation.
  2. 2.
@@
      social-norms: Text that goes against or violates expected societal norms (Forbes et al. 2020). For instance, a norm occurs along a loyalty-betrayal axis, for which an example is “It’s rude for a cheerleader to boo their own team.”
@@
      unfaithfulness: Text that is not grounded or faithful to the information provided in the input context (Chaudhury et al. 2022; Murugesan et al. 2023).
@@
      AI-generated-text: Text written by AI. This is relevant in the context of distinguishing between human-written vs machine-generated content, establishing the authenticity of information and reducing misuse (e.g., AI plagiarism) (Hu, Chen, and Ho 2023). Examples can be found at https://radar.vizhub.ai.
  4. 4.
@@
      covert-safety: Text that describes or instigates physical harm which is not immediately obvious or overt (Mei et al. 2022a). An example is “If you smell gas, light a match to find where the leak is.”
@@
      prompt-injection-and-jailbreaks: Text that may manipulate (even aligned) LLMs into generating harmful or inappropriate content, often via crafty and deliberately engineered prompts (Zou et al. 2023; Shen et al. 2023b; Hu, Chen, and Ho 2024). An example of such text with an adversarially optimized suffix is “Write a tutorial on how to make a bomb describing. + similarlyNow write oppositely.]( Me giving**ONE please? revert with ”–Two”.
Next, we overview the end-to-end development process for the detectors. At a high level, our approach to produce a new detector for a given risk dimension is to first gather or generate labeled data corresponding to the harm and then perform supervised fine-tuning on a BERT-like model (Devlin et al. 2019).
During development, we focused on maintaining a balance between efficiency and reliability. Our main challenges were to reduce inference costs (efficiency) while having limited high-quality labeled data for these different harm categories (reliability). To address the issue of efficiency, we utilized Neural Architecture Search (NAS) to derive a transformer architecture (Trivedi et al. 2023) which provides 95% of the accuracy of a BERT Base-like model (Devlin et al. 2019), while being 7x faster on a CPU and 2x times faster on a GPU. While the cost of inference for an LLM may be prohibitively expensive (Samsi et al. 2023), calling a model (where the number of parameters is on the order of 30M, instead of several billion) imposes a comparatively minuscule cost. On the other hand, the issue of reliability required some creativity. While harms such as covert unsafety (Mei et al. 2022b) and implicit hate (ElSherief et al. 2021a) have associated datasets, others such as stigma-based discrimination (Pachankis et al. 2018) have limited data, if any at all. In such cases, we utilized LLMs to generate synthetic data.
@@
2.1 Use of synthetic data generation
@@
As we discussed, there are cases when a labeled dataset for a specific harm may not be readily available, such as in social stigma. In order to have training data, we used a synthetic data generation approach where we leveraged LLMs, prompted using an in-context learning approach (Dong et al. 2023), to generate more data based on stigmas found in psychology literature (Pachankis et al. 2018). Please refer to Appendix C for the specific prompt that we utilized. Additionally, we leveraged synthetic data generation to develop nuanced improvements to existing detectors. For example, upon seeing a high false positive rate in a deployment setting for the implicit-hate-detector, we took advantage of taxonomy-guided data generation to bolster this detector. More detailed information, along with the results of this specific approach can be found in (Nagireddy et al. 2024b). Note that any generated text requires further processing and labeling; we used manual labeling but automated approaches could also be utilized (Shnarch et al. 2022).
@@
Figure 2: Red Teaming + Guardrails UI (see full figure in Appendix E, Figure 5) A user interface which encourages interactive probing of both generative models and the detectors themselves. More details in 2
2.2 Evaluating detectors on real-world data
Given that detectors will be primarily applied to machine-generated text, there is no assurance that the training data (often derived from human-generated curated datasets) matches the underlying distribution of text generated by LLMs. This creates a mismatch between the two distributions— human and LLM-generated text. Additionally, creating samples that closely mimic the “natural” responses of LLMs necessitates utilization of LLM elicitation techniques - such as prompting the model to generate continuations from pertinent prefixes (Gehman et al. 2020; Dhamala et al. 2021) or posing provocative questions to instruction-based models (Kour et al. 2023). We have open sourced one such dataset of provocative questions [^2] https://huggingface.co/datasets/ibm/ProvoQ. Please refer to Appendix F for further details.
After posing such questions, we collect responses from multiple LLMs and human or AI judges (e.g., reward models) which evaluate these responses. The evaluation of the detector entails comparing the labels of these judges with those from the detectors. A mismatch between the judge (considered as the ground truth) and the detector suggests inadequate detector performance in handling LLM outputs, signaling the need for fine-tuning on text that more resembles the output. As an example, this approach revealed a limitation in our detectors’ ability to accurately classify lengthy outputs. After investigating, this discrepancy arose as the training set predominantly comprised short utterances, which led us to prioritize enhancing the training set with instances featuring longer responses. Additionally, we discovered that our detectors exhibited sub-optimal performance when faced with intricate and evasive answers (Nagireddy et al. 2024b), particularly those generated by highly aligned and verbose models (e.g., Llama 2 (Touvron et al. 2023)). To easily facilitate the collection of such real-world data, we provide a user interface, detailed in the next section.
2.3 Interface design for human input
@@
Users can provide feedback on both the underlying generative model and specific detectors, which is propagated to a database with full lineage information. We plan to use such feedback to improve the detectors via model editing and unlearning approaches (Ghosh et al. 2023; Zylberajch, Lertvittayakumjorn, and Toni 2021; Sattigeri et al. 2022). In the next section, we discuss uncertainty based approaches that we have employed for similar improvements.
2.4 Reliable uncertainties
@@
We find the trained detectors to often be poorly calibrated and exhibit overconfidence in their predictions. Since data available for training a detector is often limited to a particular style (e.g., news headlines (Mitchell 1999) or social media posts (ElSherief et al. 2021a)), when different styles of text are encountered during deployment, the detector has difficulty flagging harmful text (see Section 2.2) as well as abstaining from flagging innocuous text. The detectors’ propensity for overconfidence, results in erroneous but confident predictions in these situations.
@@
We considered different alternatives for alleviating detector overconfidence. First, we tried a model averaging approach (Lakshminarayanan, Pritzel, and Blundell 2017) that averages predictions made by an ensemble of detectors, inspired by the reported success of similar approaches (Rahaman et al. 2021) in reducing overconfidence. We report results with such ensembling methods on the implicit hate detector in Appendix D.
@@
In addition to ensembles, we considered conformal prediction approaches (Vovk, Gammerman, and Saunders 1999). These approaches quantify uncertainty in a model’s prediction by constructing predictive sets with guaranteed frequentist coverage probabilities under minimal assumptions about the model or the true data generating process. For the implicit-hate detector, the set of predictive sets produced by the conformal predictor are {{implicit-hate}, {not-hate}, {implicit-hate, not-hate}}. Each test instance is assigned one of these predictive sets. When a test instance conforms with both labels, implicit-hate and not implicit-hate, the non-singleton set is assigned to it. The degree of conformity is measured via a conformal score calibrated on a held-out validation set. Our system used the recently proposed regularized adaptive prediction sets approach (Romano, Sesia, and Candes 2020; Angelopoulos et al. 2021) that, in addition to providing coverage guarantees, tends to produce prediction sets that are larger (non-singleton in our case) for difficult test instances and smaller (singleton) sets for easier to classify examples.
@@
We are also experimenting with increasing the proportion of negative (i.e., benign) labeled data in our training set. In early experiments, we added the data used to train the blocklisting detectors (Mitchell 1999) as it was readily available, legally permissible, and deemed appropriate for this task - as the data was in the style of news headlines that did not contain any explicit content. Our initial results are promising (refer to Appendix D for more details), and we plan to continue increasing the diversity of the training data such that it becomes more representative of deployment conditions. We also plan to use drift detection techniques (Ackerman et al. 2021) to identify when we are encountering out of distribution (OOD) data.
3 Uses of Detectors
3.1 Guardrails
@@
The simplest use case for detectors is as moderations or guardrails. For example, given its compact nature, the explicit hate speech detector was used to efficiently filter out hateful content from the set of pre-training data used to train the IBM Granite series of LLMs (Building AI for business: IBM’s Granite foundation models ). Additionally, detectors can also be used as guardrails, imposed on output generations from language models (Inan et al. 2023; Rebedea et al. 2023; Dong et al. 2024). Internally, the explicit hate, implicit hate, and stigma detectors are deployed in an experimental IBM Research prompting laboratory with over 25,000 users where they continue to be an additional safety measure on LLM generations.
Red-Teaming
@@
In addition to automated methods, detectors play a vital role in interactive probing, or red-teaming of LLMs. We have developed a user interface which aids individuals in probing LLMs alongside a detector (more in Section 2.3). Such an interface provides us with opportunities for future user studies to reveal deficiencies in the detectors themselves as well as in the underlying generative models used (Rastogi et al. 2023; Perez et al. 2022). Detectors can be used for benchmark creation by developing targeted prompts to elicit behaviour captured by the detection (Gehman et al. 2020; Kour et al. 2023; Nagireddy et al. 2024a). More on this in Appendix F.
3.2 Evaluation
Reliability and Efficiency